OpenStudy (anonymous):

Find the The domain of the function of 1/sqrt(9x+3) my answer after solving was [-2/3, infinity) When i asked for help my teacher gave me this cryptic response: "what has to be true of expressions in the denominator? Combine these ideas and your will get your answer!" I have no idea what she is talking about. Please help me.

OpenStudy (anonymous):

this is a rational function, what makes a rational function DNE? i have absolutely no idea how you arrived at [-2/3, infinity]

OpenStudy (anonymous):

9x + 3 = 0, x = -3/9 = -1/3

OpenStudy (anonymous):

because it has to be in interval notation. so the set up would be 9x+3is greater than or equal to 0, and i solved based upon that which gave me -2/3, so the interval would be [-2/3,infinity) because the number is given it has to have the [ sign for the number . the problem is actuall 1/sqrt9x+2 ...not 9x+3, i wrote it wrong on here

OpenStudy (anonymous):

What does DNE mean?